1. If the CPI for 2017 is 120 and the CPI for 2018 is 140, what is the rate of inflation between 2017 and 2018?

2. Identify and briefly describe the problems that always arise from measuring price levels with a fixed basket of goods and what steps can be taken to counter these problems.
Answer #1

1. Inflation rate = (CPI 2018 - CPI 2017 ) / CPI 2017

= (140-120)/120 = 20/120 = 1/6 = 0.167 = 16.7%.